OpenStudy (anonymous):

how do u solve (x+2)^2=58

OpenStudy (anonymous):

x+2=+-(58)^1/2 x=-2+-(58)^1/2

OpenStudy (angela210793):

x^2+4x+4-58=0 x^2+4x-54=0 x1=-2-sqr58 and x2=sqr58-2

OpenStudy (amistre64):

i would subtract out the 2 and add in a +-sqrt(58) -2 +- sqrt(58)